# Faithful Group Shapley Value

### v1

@@ Appendix A Experimental Details
A.1 Motivational Example (Section 1)
@@
The training data is partitioned into two equal-sized groups of 100 samples each. To simulate the shell company attack, the second group is further subdivided evenly into 2, 3, or 4 subgroups. This results in a total of $k\in\{2,3,4,5\}$ disjoint groups: the first group consistently contains 100 samples, while the remaining $k-1$ groups are allocated approximately $100/(k-1)$ samples each, with any remainder distributed to maintain balance (e.g., 33, 33, and 34 when $k=4$).
@@
A.2 Comparison with benchmark methods (Section 4.1)
For all methods compared here, we adjust their configurations, such that the total number of utility evaluations is aligned to the fixed number of 20,000. Note that since FGSV estimates the value for one group at a time, for fairness, the total budget is equally divided among the four groups, with 5,000 utility evaluations allocated per group. For FGSV, we set the threshold parameter to $\bar{s}=10$ and choose $m_{1}=m_{2}$ such that the total number of utility function calls sums to 20,000. Below, we describe the setup and computational structure of each method in detail, largely based on the implementation and descriptions from [24].
  • Permutation estimator [2, 6].
@@
  • Group Testing estimator [12, 34].
@@
  • Complement Contribution estimator [38].
@@
    This expression enables the estimator to reuse utility evaluations for both $S$ and its complement $[n]\setminus S$. At each iteration, it samples a subset size $s\in[n]$ uniformly and then uniformly samples a subset $S\subseteq[n]$ of size $s$. Then, the final estimator is:
@@
  • One-for-All estimator [24].
    The One-for-All estimator is based on the following alternative formulation of the Shapley value [^1] While the One-for-All estimator can estimate a variety of values (e.g., Banzhaf [33], Beta-Shapley [19]), we focus on its use for Shapley value estimation.:
@@
    This formulation allows each sampled subset to contribute to the estimates of all players, enabling efficient sample reuse. First, it deterministically allocates $22n+$ 2 utility evaluations for subset sizes $s\in\{0,1,n-1,n\}$, and uses them to compute the corresponding expectations exactly. Then, for the remaining possible subset sizes $\{2,\dots,n-2\}$, it samples $s$ with a predefined sampling probability $q(s)$, and then samples a subset $S_{t}\subseteq[n]$ of size $s$ uniformly at random. The final estimation is:
@@
    where $T_{i,s}^{\text{in}}$ and $T_{i,s}^{\text{out}}$ denote the number of Monte Carlo samples of size $s$ in which $i$ is included and excluded, respectively. In our experiment, we used $q(s)\propto\frac{1}{\sqrt{s(n-s)}}$, which is proved to be the optimal sampling distribution for Shapley value estimation. Next, after allocating $22n+$ 2 evaluations deterministically for subset sizes $s\in\{0,1,n-1,n\}$, the remaining evaluations are used for Monte Carlo sampling. That is, we set $T=20000-(2n+2)$ for sampling subset sizes $s\in\{2,\dots,n-2\}$.
  • KernelSHAP [26].
@@
Figure 5: Empirical average runtime (in seconds) of $U(S)$ evaluation as a function of subset size $s=|S|$. Each curve represents the mean over 50 randomly sampled subsets of size $s$; shaded areas indicate ±$\pm 1$ standard deviation.
@@
A.3 Faithful copyright attribution in generative AI (Section 4.2)
We evaluate SRS and FSRS in the context of group-level copyright attribution for generative models, using a logo generation task based on Stable Diffusion. To facilitate comparison, we mostly followed the experimental setup in [35]. Nonetheless, for clarity and completeness, we summarize the relevant details below.
To construct the experiment, we select four logo classes (Google, Sprite, Vodafone, and Starbucks) for evaluation, and use the remaining 23 brands to initialize a baseline model via fine-tuning of Stable Diffusion v1.4 [28]. We then assess the impact of fine-tuning with each of the four excluded brands by measuring how their inclusion alters the generation behavior of the model.
Fine-tuning is performed using Low-Rank Adaptation (LoRA) [11], a technique that inserts trainable low-rank matrices into the attention layers of the diffusion model. The rank $r$ and scaling factor $\alpha$ are both set to 8, and fine-tuning is only applied to the attention layers of the UNet module, while all other weights of the Stable Diffusion model are kept frozen. This configuration ensures efficient and stable adaptation to the small-scale dataset, in line with standard LoRA usage. [^2] https://huggingface.co/blog/lora
@@
Following the formulation of [35], we define the utility function as the log-likelihood of generated images. For a given length of the denoising steps $T$, let $(x_{T},\dots,x_{1},x_{0})$ denote the reverse trajectory defined by the diffusion scheduler, where $x_{T}\sim\mathcal{N}(0,I)$ is the initial latent noise, and $x_{0}=x_{(\text{gen})}$ is the final generated image. The likelihood of $x_{(\text{gen})}$ under model parameters $\theta$ is defined as:
@@
A.4 Faithful explainable AI (Section 4.3)
We evaluated FGSV and GSV on the Diabetes dataset [5], which includes 442 samples with 10 demographic and health-related features. The dataset is preprocessed by centering and standardizing the response variable and splitting the dataset into 350 training and 92 test samples using a fixed random seed.
@@
Appendix B Approximation algorithm with non-informative data augmentation
@@
3:  for $s=1$ to $n-1$ do
4:     Set $s_{1}^{*}\leftarrow\lfloor s\alpha_{0}\rfloor$. {Expected intersection size for current $s$}

### v2

@@ Appendix B Experimental Details
B.1 Motivational Example (Section 1)
@@
The training data is partitioned into two equal-sized groups of 100 samples each. To simulate the shell company attack, the second group is further subdivided evenly into 2, 3, or 4 subgroups. This results in a total of $k\in\{2,3,4,5\}$ disjoint groups: the first group consistently contains 100 samples, while the remaining k1k-1 groups are allocated approximately 100/(k1)100/(k-1) samples each, with any remainder distributed to maintain balance (e.g., 33, 33, and 34 when $k=4$).
@@
B.2 Comparison with benchmark methods (Section 4.1)
For all methods compared here, we adjust their configurations, such that the total number of utility evaluations is aligned to the fixed number of 20,000. Note that since FGSV estimates the value for one group at a time, for fairness, the total budget is equally divided among the four groups, with 5,000 utility evaluations allocated per group. For FGSV, we set the threshold parameter to $\bar{s}=10$ and choose $m_{1}=m_{2}$ such that the total number of utility function calls sums to 20,000. Below, we describe the setup and computational structure of each method in detail, largely based on the implementation and descriptions from Li and Yu [2024b].
  • Permutation estimator [Castro et al., 2009, Ghorbani and Zou, 2019].
@@
  • Group Testing estimator [Jia et al., 2019, Wang and Jia, 2023a].
@@
  • Complement Contribution estimator [Zhang et al., 2023].
@@
    This expression enables the estimator to reuse utility evaluations for both $S$ and its complement $n]\setminus S$. At each iteration, it samples a subset size $s\in[n]$ uniformly and then uniformly samples a subset $S\subseteq[n]$ of size $s$. Then, the final estimator is:
@@
  • One-for-All estimator [Li and Yu, 2024b].
    The One-for-All estimator is based on the following alternative formulation of the Shapley value [^1] While the One-for-All estimator can estimate a variety of values (e.g., Banzhaf [Wang and Jia, 2023b], Beta-Shapley [Kwon and Zou, 2022]), we focus on its use for Shapley value estimation.:
@@
    This formulation allows each sampled subset to contribute to the estimates of all players, enabling efficient sample reuse. First, it deterministically allocates $2n+2$ utility evaluations for subset sizes $s\in\{0,1,n-1,n\}$, and uses them to compute the corresponding expectations exactly. Then, for the remaining possible subset sizes $\{2,\dots,n-2\}$, it samples $s$ with a predefined sampling probability $q(s)$, and then samples a subset $S_{t}\subseteq[n]$ of size $s$ uniformly at random. The final estimation is:
@@
    where $T_{i,s}^{\text{in}}$ and $T_{i,s}^{\text{out}}$ denote the number of Monte Carlo samples of size $s$ in which $i$ is included and excluded, respectively. In our experiment, we used $q(s)\propto\frac{1}{\sqrt{s(n-s)}}$, which is proved to be the optimal sampling distribution for Shapley value estimation. Next, after allocating $2n+2$ evaluations deterministically for subset sizes $s\in\{0,1,n-1,n\}$, the remaining evaluations are used for Monte Carlo sampling. That is, we set T=20000(2n+2)T=20000-(2n+2) for sampling subset sizes $s\in\{2,\dots,n-2\}$.
  • KernelSHAP [Lundberg, 2017].
@@
Figure 5: Empirical average runtime (in seconds) of $U(S)$ evaluation as a function of subset size $s=|S|$. Each curve represents the mean over 50 randomly sampled subsets of size $s$; shaded areas indicate $\pm 1$ standard deviation.
@@
B.3 Faithful copyright attribution in generative AI (Section 4.2)
We evaluate SRS and FSRS in the context of group-level copyright attribution for generative models, using a logo generation task based on Stable Diffusion. To facilitate comparison, we mostly followed the experimental setup in Wang et al. [2024]. Nonetheless, for clarity and completeness, we summarize the relevant details below.
To construct the experiment, we select four logo classes (Google, Sprite, Vodafone, and Starbucks) for evaluation, and use the remaining 23 brands to initialize a baseline model via fine-tuning of Stable Diffusion v1.4 [Rombach et al., 2022]. We then assess the impact of fine-tuning with each of the four excluded brands by measuring how their inclusion alters the generation behavior of the model.
Fine-tuning is performed using Low-Rank Adaptation (LoRA) [Hu et al., 2022], a technique that inserts trainable low-rank matrices into the attention layers of the diffusion model. The rank $r$ and scaling factor $\alpha$ are both set to 8, and fine-tuning is only applied to the attention layers of the UNet module, while all other weights of the Stable Diffusion model are kept frozen. This configuration ensures efficient and stable adaptation to the small-scale dataset, in line with standard LoRA usage. [^2] https://huggingface.co/blog/lora
@@
Following the formulation of Wang et al. [2024], we define the utility function as the log-likelihood of generated images. For a given length of the denoising steps $T$, let $(x_{T},\dots,x_{1},x_{0})$ denote the reverse trajectory defined by the diffusion scheduler, where $x_{T}\sim\mathcal{N}(0,I)$ is the initial latent noise, and $x_{0}=x_{(\text{gen})}$ is the final generated image. The likelihood of $x_{(\text{gen})}$ under model parameters $\theta$ is defined as:
@@
B.4 Faithful explainable AI (Section 4.3)
We evaluated FGSV and GSV on the Diabetes dataset [Efron et al., 2004], which includes 442 samples with 10 demographic and health-related features. The dataset is preprocessed by centering and standardizing the response variable and splitting the dataset into 350 training and 92 test samples using a fixed random seed.
@@
Appendix C Approximation algorithm with non-informative data augmentation
@@
3:for $s=1$ to n1n-1 do
4:  Set $s_{1}^{*}\leftarrow\lfloor s\alpha_{0}\rfloor$. {Expected intersection size for current ss}